# Safe, Task-Consistent Manipulation with Operational Space Control Barrier Functions

### v1

@@ Abstract
Safe real-time control of robotic manipulators in unstructured environments requires handling numerous safety constraints without compromising task performance. Traditional approaches, such as artificial potential fields (APFs), suffer from local minima, oscillations, and limited scalability, while model predictive control (MPC) can be computationally expensive. Control barrier functions (CBFs) offer a promising alternative due to their high level of robustness and low computational cost, but these safety filters must be carefully designed to avoid significant reductions in the overall performance of the manipulator. In this work, we introduce an Operational Space Control Barrier Function (OSCBF) framework that integrates safety constraints while preserving task-consistent behavior. Our approach scales to hundreds of simultaneous constraints while retaining real-time control rates, ensuring collision avoidance, singularity prevention, and workspace containment even in highly cluttered and dynamic settings. By explicitly accounting for the task hierarchy in the CBF objective, we prevent degraded performance across both joint-space and operational-space tasks, when at the limit of safety. Our open-source, high-performance software will be available at our project webpage, https://stanfordasl.github.io/oscbf/
Figure 1: Safe and performant highly-constrained manipulation. Our OSCBF controller maintains safety for hundreds of constraints enforced concurrently, using the full second-order robot dynamics and torque control, and operates at real-time control rates (over 1000 Hz). Shown above is the simultaneous constraint evolution for all CBFs during one adversarial trajectory from teleoperation. Across 168 safety conditions, the robot remains safe ($h(\mathbf{z})>0$) without over-conservative behavior near the boundary of safety ($h(\mathbf{z})=0$).
@@ I-A Statement of Contributions
We present a framework for integrating Control Barrier Functions (CBFs) into Operational Space Contol (OSC) which (1) enforces safety while respecting the hierarchical task definition of OSC across the operational and joint spaces, (2) applies to both kinematically- and dynamically- controlled manipulators, and (3) scales to large numbers of safety conditions while maintaining real-time control rates, even when performing dynamic motions through highly-occluded environments. Additionally, we release our software, CBFpy [18], which provides an easy-to-use and high-performance interface for CBF controller design. The code, along with additional media, will be made available at https://stanfordasl.github.io/oscbf/
@@ I-B Paper Organization
In Section II, we review the construction of control barrier functions for constraints of relative degree 1 and 2, and discuss practical application of CBFs for highly-constrained systems. In Section III, we review manipulator kinematics, dynamics, and tasks. In Section IV, we detail the necessity of task-consistent safety filters, then we construct our Operational Space Control Barrier Function (OSCBF) controller, for both velocity- and torque- controlled robots. In Section V, we present analysis of the OSCBF controller’s performance and computational efficiency for five common types of safety constraints, then, in simulation, we evaluate the controller in highly constrained environments and during dynamic motions.
@@ IV-A Task Consistency
When imposing a safety constraint on a robotic manipulator, many trajectories can maintain safety, but not all can additionally maximize task performance. As shown in Fig. 2, while “task-inconsistent” CBFs can remain safe (i.e., $h(\mathbf{z})\geq 0$), they can also introduce excess motion at the boundary of safety. Examples of “task-inconsistency” include: (1) optimizing a joint-space metric, while the task is defined in the operational space; (2) the converse: optimizing an operational-space metric, while ignoring a secondary null space joint task; (3) optimizing over torques, rather than accelerations.
In general, task consistency implies that rather than applying the CBF safety filter directly to the control input, the filter should minimally modify an output that reflects the task and hierarchy definition. By doing so, this eliminates unnecessary motion when the reference command moves further into the unsafe set. Inconsistent examples (1) and (2) above misrepresent the task hierarchy. (3) optimizes the input (torque), which less directly reflects the position-based tasks than acceleration, due to the additional inertial mapping involved.
@@ V-A Implementation
Experiment CBF Constraints Velocity Control Torque Control Torque Control (Dynamic)
Mean Std. Mean Std. Mean Std.
Singularity avoidance 1 10.73 0.84 7.22 0.35 2.95 0.09
End-effector safe-set containment 6 22.49 1.46 15.12 0.92 3.54 0.12
Joint limit avoidance 14 13.41 0.83 12.13 1.16 3.36 0.12
Whole-body collision avoidance 21 5.92 1.31 4.92 0.94 2.60 0.37
Whole-body safe-set containment 126 4.96 0.37 4.14 0.25 2.24 0.08
(All of the above) 168 3.21 0.59 3.13 0.57 1.89 0.34
@@
The Python interface is sufficiently fast for kilohertz control rates (see Table II), even when the full manipulator dynamics are evaluated, and with over 100 CBF constraints. The centrifugal/Coriolis forces tend to be somewhat expensive to compute, and if the robot motion is not highly dynamic, these can be neglected for more efficient computation. If real-time C++ code is required, the OSCBF can be directly converted to a high-level optimized XLA representation and called via the XLA C++ interface. Compute times are from an Intel i7-1360p (5 GHz) NUC computer with 64 GB RAM, and Jax v.0.4.30. JIT compilation times are typically 2-5 seconds.
@@ V-B4 Whole-body collision avoidance
for all $i$ positions $\mathbf{x}_{p,i}(\mathbf{q})$ and radii $r_{i}$ in the collision model of the robot, after propagation through the forward kinematics, and all $j$ positions/radii of the environmental collision model, $\mathbf{x}_{p,j,\text{obs}}$ and $r_{j,\text{obs}}$. Our simple collision model for the Franka Panda comprises 21 spheres, and we emphasize that this will still work with a more detailed collision model.
@@ V-C Highly-Constrained Environments
Figure 4: Scaling up collision avoidance. Even in highly-cluttered scenes, our OSCBF controller maintains safety, task-tracking performance, and real-time control rates. Consider a tabletop environment (left) with many randomly-generated collision bodies, shown in blue. OSCBF scales to over 400 CBF constraints while retaining real-time control rates (1000 Hz) for torque control, and well over 1000 constraints while retaining good control rates (100 Hz) for velocity control. Here, (Dynamic) indicates compensation for centrifugal and Coriolis forces under high-speed motions, and we indicate both the mean frequency and the minimum frequency, assuming an allowable 5% packet drop.
@@ V-D Dynamic Safety
Manipulators are often capable of high-speed dynamic motions, but guaranteeing safety during these motions can be difficult, particularly in an online setting without pre-planned trajectories.
Previous works assumed that a low-level tracking controller can guarantee that a safe velocity command can always be met, but we emphasize that this is not the case for dynamic motions, with torque limits. As shown in Fig. 5, making this assumption can sometimes maintain safety (here, the end-effector trajectory does remain within the safe set), but the desired safe velocity reference cannot always be met, given the torque limits. This leads to a degradation in task tracking performance, as indicated by the departure from the desired straight-line motion. However, if the full second-order dynamics of the manipulator are considered, with explicit constraints on the torque input, we can maintain both tracking performance and safety.
In general, if the robot control interface allows for torque inputs, this should be used over a velocity control interface if both safety and dynamic motions are needed. If only velocity control is available, highly dynamic motions should be avoided. Remark: Further analysis on sending velocity commands from an integrated torque OSCBF may be warranted.
V-E Discussion
V-E1 Parameter tuning
@@ VI Conclusion
Future work will focus on hardware validation, for both velocity- and torque-controlled robots, and mobile manipulators. Moving forward, we aim to apply this controller as a core part of imitation-learning-based manipulation policy training and deployment: collecting data safely via teleoperation, and deploying the learned policies with this safety filter maintaining whole-body safety of the robot.

### v2

@@ Abstract
Safe real-time control of robotic manipulators in unstructured environments requires handling numerous safety constraints without compromising task performance. Traditional approaches, such as artificial potential fields (APFs), suffer from local minima, oscillations, and limited scalability, while model predictive control (MPC) can be computationally expensive. Control barrier functions (CBFs) offer a promising alternative due to their high level of robustness and low computational cost, but these safety filters must be carefully designed to avoid significant reductions in the overall performance of the manipulator. In this work, we introduce an Operational Space Control Barrier Function (OSCBF) framework that integrates safety constraints while preserving task-consistent behavior. Our approach scales to hundreds of simultaneous constraints while retaining real-time control rates, ensuring collision avoidance, singularity prevention, and workspace containment even in highly cluttered settings or during dynamic motions. By explicitly accounting for the task hierarchy in the CBF objective, we prevent degraded performance across both joint-space and operational-space tasks, when at the limit of safety. We validate performance in both simulation and hardware, and release our open-source high-performance code and media on our project webpage, https://stanfordasl.github.io/oscbf/
Figure 1: Safe and performant highly-constrained manipulation. Our OSCBF controller maintains safety for hundreds of constraints enforced concurrently, using the full second-order robot dynamics and torque control, and operates at real-time control rates (over 1000 Hz). Shown above is the simultaneous constraint evolution for all CBFs during one adversarial trajectory from teleoperation (end-effector target shown in red). Across 168 safety conditions, the robot remains safe ($h(\mathbf{z})>0$) without over-conservative behavior near the boundary of safety ($h(\mathbf{z})=0$).
@@ I-A Statement of Contributions
We present a framework for integrating Control Barrier Functions (CBFs) into Operational Space Contol (OSC) which (1) enforces safety while respecting the hierarchical task definition of OSC across the operational and joint spaces, (2) applies to both kinematically- and dynamically- controlled manipulators, and (3) scales to large numbers of safety conditions while maintaining real-time control rates, even when performing dynamic motions through highly-occluded environments. Additionally, we release our software, CBFpy [18], which provides an easy-to-use and high-performance interface for CBF controller design. The code, along with additional media, is available at https://stanfordasl.github.io/oscbf/
@@ I-B Paper Organization
In Section II, we review the construction of control barrier functions for constraints of relative degree 1 and 2, and discuss practical application of CBFs for highly-constrained systems. In Section III, we review manipulator kinematics, dynamics, and tasks. In Section IV, we detail the necessity of task-consistent safety filters, then we construct our Operational Space Control Barrier Function (OSCBF) controller, for both velocity- and torque- controlled robots. In Section V, we present analysis of the OSCBF controller’s performance and computational efficiency for five common types of safety constraints, then, in simulation and on hardware, we evaluate the controller in highly constrained environments and during dynamic motions.
@@ IV-A Task Consistency
When imposing a safety constraint on a robotic manipulator, many trajectories can maintain safety, but not all can additionally maximize task performance. As shown in Fig. 2, while “task-inconsistent” CBFs can remain safe (i.e., $h(\mathbf{z})\geq 0$), they can also introduce excess motion at the boundary of safety. Examples of “task-inconsistency” include: (1) optimizing a joint-space metric, while the task is defined in the operational space; (2) the converse: optimizing an operational-space metric, while ignoring a secondary null space joint task; (3) optimizing a torque-based metric, rather than an acceleration-based metric.
In general, task consistency implies that rather than applying the CBF safety filter directly to the control input, the filter should minimally modify an output that reflects the task and hierarchy definition. By doing so, this eliminates unnecessary motion when the reference command moves further into the unsafe set. Inconsistent examples (1) and (2) above misrepresent the task hierarchy. (3) optimizes a metric which less directly reflects the position-based tasks than acceleration, due to the additional inertial mapping involved.
@@ V-A Implementation
Experiment CBF Constraints Velocity Control Torque Control
Mean Lower bound (5th percentile) Mean Lower bound (5th percentile)
Singularity avoidance 1 9.55 5.15 6.14 3.89
End-effector safe-set containment 6 9.49 6.19 7.17 4.12
Joint limit avoidance 14 9.99 6.15 7.72 4.45
Whole-body collision avoidance 21 7.48 4.40 5.71 3.06
Whole-body safe-set containment 126 4.94 3.33 3.89 2.59
(All of the above) 168 3.24 2.35 2.94 2.25
@@
The Python interface is sufficiently fast for kilohertz control rates (see Table II), even when the full manipulator dynamics are evaluated, and with over 100 CBF constraints. If C++ code is required, the OSCBF can be directly converted to a high-level optimized XLA representation and called via the XLA C++ interface. Compute times are from an Intel i7-1360p (5 GHz) NUC computer with 64 GB RAM, and Jax v.0.4.30. JIT compilation times are typically 2-5 seconds.
@@ V-B4 Whole-body collision avoidance
for all $i$ positions $\mathbf{x}_{p,i}(\mathbf{q})$ and radii $r_{i}$ in the collision model of the robot, after propagation through the forward kinematics, and all $j$ positions/radii of the environmental collision model, $\mathbf{x}_{p,j,\text{obs}}$ and $r_{j,\text{obs}}$. Our simple collision model for the Franka Panda consists of 21 spheres, and we emphasize that this will still work with a more detailed collision model.
@@ V-C Highly-Constrained Environments
Figure 4: Scaling up collision avoidance. Even in highly-cluttered scenes, our OSCBF controller maintains safety, task-tracking performance, and real-time control rates. Consider a tabletop environment (inset) with many randomly-generated collision bodies, shown in blue. OSCBF scales to over 400 CBF constraints while retaining real-time control rates (1000 Hz) for torque control, and well over 1000 constraints while retaining good control rates (100 Hz) for velocity control. We indicate both the mean frequency and the minimum frequency, assuming an allowable 5% packet drop.
@@ V-D Dynamic Safety
Manipulators are often capable of high-speed dynamic motions, but guaranteeing safety during these motions can be difficult, particularly in an online setting without pre-planned trajectories. Previous works assumed that a low-level tracking controller can guarantee that a safe velocity command can always be met, but we emphasize that this is not the case for dynamic motions, with torque limits. As shown in Fig. 5, making this assumption can sometimes maintain safety (here, the end-effector trajectory does remain within the safe set), but the desired safe velocity reference cannot always be met, given the torque limits. This leads to a degradation in task tracking performance, as indicated by the departure from the desired straight-line motion. However, if the full second-order dynamics of the manipulator are considered, with explicit constraints on the torque input, we can maintain both tracking performance and safety.
V-E Hardware Experiments
To validate OSCBF’s performance on hardware, we set up a 1 kHz real-time torque control interface for safe teleoperation and operational space control of the Franka Panda. Across four experiments designed to replicate the simulation results, OSCBF demonstrates remarkably smooth, low-latency tracking of the end-effector commands, safely approaching the boundary of safety within just 6 mm (Fig. 6 A/B/D) or quickly moving in and out of near-singular configurations ($\mu=1\mathrm{e}{-2}$) (Fig. 6 C) while operating at the dynamic limits of the Franka hardware. Fig. 6 B also demonstrates whole-body collision avoidance while maintaining good tracking of the operational space task, by filtering the secondary null space posture task when constrained by the environment. Notably, we observe effectively zero mismatch between the hardware data and simulation, save for a miniscule amount of latency (50 ms) from the teleoperation interface. Videos of the experiments can be found on our project webpage.
Figure 6: Hardware validation. We evaluate OSCBF’s performance across four scenarios on the Franka Emika Panda hardware. In three teleoperated experiments, we demonstrate (A) whole-body workspace containment, (B) whole-body collision avoidance, and (C) singularity avoidance. In (D), we command a fast sinusoidal end-effector trajectory which, without OSCBF, would lead directly to a collision. The plot visualizes the minimum CBF constraint evolution in this experiment, indicating smooth, safe, and non-overly-conservative behavior from the OSCBF controller.
V-F Discussion
V-F1 Parameter tuning
@@ VI Conclusion
Future work will focus on applications on high-DOF mobile manipulators or bimanual systems, and integration with real-time perception. Moving forward, we aim to apply this controller as a core part of imitation-learning-based manipulation policy training and deployment: collecting data safely via teleoperation, and deploying the learned policies with OSCBF maintaining whole-body safety of the robot.